# From Wavefunctional Entanglement to Entangled Wavefunctional Degrees of Freedom

### v1

@@ Abstract
Conclusions.—To summarize, I have examined the equivalence of entanglement between optical modes and entanglement between optical photons that are described by such modes, and have concluded that it is, indeed, possible to transform the former into the latter; usually, the latter is harder to realize and is more useful than the former. Specifically, I have shown that systems having four or more spatial modes, and two or more input unentangled photons—such as non-Abelian, quantum holonomic systems [89, 90, 91] in which light field modes are in photon number entangled relationships with each other [74]—can be used to nonlocally entangle photons. Highly entangled optical systems can be used to construct modular, programmable platforms that are more readily interfaced and integrated with existent long-distance quantum communication networks, which already employ optical photons encoded as flying qubits. Additionally, a future avenue of fundamental research might entail placing the results of this Letter on a firmer mathematical grounding, and extending them to quantum systems found in other branches of physics.
Note added.—Shortly before submitting this paper, I became aware of a recently posted arXiv preprint [92], by the Zeilinger group in Vienna, that sheds new light on the subjective nature of path information, specifically, the “which-way information” in three-particle interferometry. This path information plays a key role in how spatial, “mode-entangled,” or path-entangled superpositions of photons are described.
@@ Acknowledgements.
Acknowledgments.—I would like to thank Chandra Raman for illuminating discussions on the differences between entangled modes and entangled particles, as well as for kindly providing key references on this topic.Author Contributions.—A.B. conceptualized this work, carried out all the calculations and derivations, prepared all the figures, and wrote the manuscript.
Data Availability.—This publication reports a theoretical work, and consequently, no experimental research data were created or analyzed, during the course of this research. In particular, the author confirms that all the relevant information—namely, theoretical calculations, derivations, and accompanying results—that support the findings of this work are available within the article, and its end matter and supplementary material.
Appendix A End Matter
Entanglement measurement contexts in real and Hilbert spaces.—To distinguish between the contextual natures of photon-photon and mode-mode entanglement, consider a quantum optical re-interpretation of the E.P.R.B. Gedankenexperiment, as shown in Fig. 3(a). If we regard the horizontally and vertically polarized, single-photon states, and as the basis states—for example, and —then the corresponding, unitarily transformed states at the measurement station A (orientation angle, ; axis of linear analyzer, ) are:

### v2

@@ VII Conclusions and Outlook
In this paper, I have examined the equivalence of entanglement between optical modes and entanglement between optical photons that are described by such modes, and have concluded that it is, indeed, possible to transform the former into the latter; usually, the latter is harder to realize and is more useful than the former. Specifically, I have shown that systems having four or more spatial modes, and two or more input unentangled photons—such as non-Abelian, quantum holonomic systems [111, 112, 113] in which light field modes are in photon number entangled relationships with each other [84]—can be used to nonlocally entangle photons. This theoretical demonstration might have implications for crafting protocols and devising algorithms—that, for instance, start with mode-mode entangled states, which are, subsequently, converted to particle-particle entangled states, during a key, intermediate step—with continuous wave, quantum optical systems.
Highly entangled optical systems can be used to construct modular, programmable platforms that are more readily interfaced and integrated with existent long-distance quantum communication networks, which already employ optical photons encoded as flying qubits. Additionally, a future avenue of fundamental research might entail placing the results of this paper on a firmer mathematical grounding, and extending them to quantum systems found in other branches of physics.
@@ Acknowledgements.
I would like to thank Chandra Raman for illuminating discussions on the differences between entangled modes and entangled particles, as well as for kindly providing key references on this topic.VIII Data Availability
This paper reports a theoretical work, and consequently, no experimental research data were created or analyzed, during the course of this research. In particular, the author confirms that all the relevant information—namely, theoretical calculations, derivations, and accompanying results—that support the findings of this work are available within the article.
Appendix A Entanglement Measurement Contexts in Real and Hilbert Spaces.
To distinguish between the contextual natures of photon-photon and mode-mode entanglement, consider a quantum optical re-interpretation of the E.P.R.B. Gedankenexperiment, as shown in Fig. 3(a). If we regard the horizontally and vertically polarized, single-photon states, and as the basis states—for example, and —then the corresponding, unitarily transformed states at the measurement station A (orientation angle, ; axis of linear analyzer, ) are:
@@
Appendix B A Physical Interpretation of the Unitarily Transformed State
Remarkably, the state in Eq. 1, provides a direct and manifestly nonlocal description of the entangled particle pair throughout the flight of the photons. I am assuming that this description is based on, for example, knowledge of successful initial state preparation. In contrast, in the absence of such knowledge, one could interpret the state in Eq. 2, as encoding reconstructed descriptions of quantum events—or assignments of state vectors—after the quantum measurements have been made; notably, similar ideas have been proposed to explain the observation of delayed-choice entanglement swapping [114, 115] in which photons become entangled, after they have already been registered.
More specifically—for the case of photon-photon entanglement—the observers at the two measurement stations, A and B [see, for example, Fig. 3(a)] would assign—after performing local measurements on their individual quantum subsystems, and subsequently comparing their expectation catalogs of single and joint detection events—such ex post facto, or after-the-event quantum state assignments to describe their observations. Equation 8 gives the probabilities of all such single and joint detection events.
To properly interpret , one should regard the individual, recorded events—for instance, measurements that reveal probability amplitudes, as well as relative phases at the two space-like separated, measurement stations—as more fundamental than the quantum state, itself. In this viewpoint, therefore, the wavefunction is the description that the observers assign to the overall situation or the entire phenomenon—as it were, after all the events have occurred—and is determined by the complete set of all possible experimental arrangements; for example, of all allowable relative orientations of and in Fig. 3(a). Of note, this conception of is in harmony with several aspects of the Copenhagen interpretation of quantum mechanics [116], especially, as emphasized by Zeilinger and co-authors in the past few decades [12, 117, 13, 14]. The interpretation of the results of a recent elegant experiment, clarifying the subjective nature of path information, specifically, the “which-way information” in three-particle interferometry, also suggests the primacy of quantum events [118]. This path information plays a key role in how spatial, “mode-entangled,” or path-entangled superpositions of photons are described.
As is shown in Fig. 3(b), the degrees of entanglement of and are identical, regardless of the relative phase, . This invariance of the value of the entanglement with regard to the relative orientation (or basis) is a characteristic defining property of entangled particles. An interesting consequence of this feature is described below.
Recently, experimental violation of Bell’s inequalities due to quantum indistinguishability by path identity, as opposed to quantum entanglement, has been reported [119]. As the authors themselves emphasize, to show this effect, they were actively manipulating the state during its creation, rather than merely measuring the properties of an unaltered entangled state [119]. This manipulation is afforded by their experimental setup. In contrast, the usual setup [see, for example, Fig. 3(a)] does not change while it varies for the purposes of demonstrating the violation of Bell’s inequalities. Therefore, in the usual case, the unchanging entangled state, , or, equivalently, is responsible for the violation of Bell’s inequalities, and the origin of Bell nonlocality.
Appendix C A More Formal Justification for the Equivalence
It is well known that the modes, in quantum mechanics, can be mathematically treated as quantum states. In our approach, we model the spatial optical modes, just like any other bosonic mode, as single-particle wavefunctions. Specifically, we map such modes onto the single-quantum, energy eigenmodes of a simple harmonic oscillator. For example, to model the key term of the input state [see Eq. 3], that is, which corresponds to two distinguishable photons occupying two distinct modes, we make the following mappings:
where and are the single-particle, harmonic oscillator energy eigenfunctions, and and represent the coordinates of the two particles, respectively. Therefore, we can represent this key term as: . Strictly speaking, one could model the two optical modes more rigorously as energy eigenmodes of two harmonic oscillators in two distinct Hilbert spaces, respectively; however, one would arrive at exactly the same result. For simplicity, and without loss of generality, I am assuming that we have a single harmonic potential in real space. Let us also consider a global and time-independent anharmonic perturbation that can be applied globally, and that alters the single-particle wavefunctions, from and , to and , respectively in a smooth and continuous fashion. As is well known, such energy eigenfunctions of the anharmonically perturbed potential exist, and can be calculated using standard, time-independent perturbation methods.
Figure 2 describes a simple mechanism that involves making a quantum measurement on an ancillary photon, and that uses a quantum mechanical random process to control the degree of anharmonicity of the potential. Since the successful photo-detection of the incident photon, which is propagating through the Young’s double-slit apparatus and is interfering with itself, is a quantum mechanically probabilistic process, the state of the entire photo-detecting apparatus is brought into a macroscopic superposition of having detected and having not detected the photon. Consequently, the initial state, is transformed into this entangled superposition: , where and are suitable normalizing, expansion coefficients.
Now, we can make these inverse mappings to recover the optical modes from the corresponding energy eigenfunctions of the anharmonically perturbed potential:
where the spatial coordinates, and , and the particle labels, and have the same meanings, as before. Notice that the above step is the converse of the one shown in Eq. 15, and assumes that these mappings are invertible, which is usually the case. For simplicity, and to avoid any inadvertent pathologies, I am going to assume that this mapping is one-to-one and onto.
Therefore, the initial state, has been transformed to:
which is a genuinely entangled state of two interacting particles. Of note, as is evident from this simple argument, creating such a state requires access to at least four distinct spatial optical modes, and two incident and initially unentangled optical photons. Notice that this transformation process is not completely deterministic, as it is conditioned on the probabilistic photo-detection of the ancillary, trigger photon 3.